# Designing Library of Skill-Agents for Hardware-Level Reusability

### 04 Mar 2024 v1

@@ Abstract
To use new robot hardware in a new environment, it is necessary to develop a control program tailored to that specific robot in that environment. Considering the reusability of software among robots is crucial to minimize the effort involved in this process and maximize software reuse across different robots in different environments. Although recent generative AI has made it possible to automatically generate some software, robot programs are still difficult to generate because of robot’s physical interaction with the environment. This paper proposes a method to remedy this process by considering hardware-level reusability, using Learning-from-observation (LfO) paradigm with a pre-designed skill-agent library. The LfO framework represents the required actions in hardware-independent representations, referred to as task models, from observing human demonstrations, capturing the necessary parameters for the interaction between the environment and the robot (Ikeuchi et al. (2021)). When executing the desired actions from the task models, a set of skill agents is employed to convert the representations into robot commands. This paper focuses on the latter part of the LfO framework, utilizing the set to generate robot actions from the task models, and explores a hardware-independent design approach for these skill agents. These skill agents are described in a hardware-independent manner, considering the relative relationship between the robot’s hand position and the environment. As a result, it is possible to execute these actions on robots with different hardware configurations by simply swapping the inverse kinematics solver. This paper, first, defines a necessary and sufficient skill-agent set corresponding to cover all possible actions, and considers the design principles for these skill agents in the library. We provide concrete examples of such skill agents and demonstrate the practicality of using these skill agents by showing that the same representations can be executed on two different robots, Nextage and Fetch, using the proposed skill-agents set.
@@ 1 Introduction
Robot developers develop various types of robots, such as bipedal robots and mobile manipulators for satisfying users’ various demands. If we consider manipulation aspects, there are various types: a single-arm robot or a dual-arm robot and the degrees of freedom (DOF) of an arm is from 5 (e.g., HSR, Toyota) to 7 (e.g., Fetch Mobile Manipulator, Fetch Robotics). Some robots have additional DOF on their waist.
@@
We have been developing a robot system based on a Learning-from-observation (LfO) framework in Ikeuchi et al. (2021) [^1] In the machine-learning (ML) community, Learning-from-observation (LfO) is used with a slightly different definition. However, following the rationale mentioned in the introduction of the previous paper (Ikeuchi et al. (2021)), LfO is specifically referred to as a method that transforms input into symbolic representations based on top-down knowledge and subsequently maps them to robot actions.. Our goal is to have a robot reproduce the target behavior by simply demonstrating it in front of the robot. Thus, even users without robotics knowledge can have the robot perform tasks that they desire. Unlike similar frameworks, such as learning-from-demonstration and imitation learning (Schaal (1999); Schaal et al. (2003); Billard et al. (2008); Asfour et al. (2008); Dillmann et al. (2010); Akgun et al. (2012)), the LfO system preforms indirect behavior mimicry by first a task-encoding step converting the demonstration into an abstract intermediate representation, referred to as a task model, and then a task-decoding step converting this intermediate representation into the behavior of each robot. The task-encoding recognizes behavior as symbolic task sequences (such as a sequence of pick up, place, and release actions) and then extracts from the demonstration the parameters that are pre-defined in each task (such as where to grasp and where to put). Because only task-specific information is extracted from the demonstration, unnecessary parts are ignored, allowing the demonstrator to focus only on the important demonstration parts that need to be taught. For task-decoding, we prepare in advance the skill agents, which correspond to symbolic actions, following the design of the task model. Then, the reproduction by a robot is realized by calling the corresponding skill agent with the observed parameters.
This paper focuses on the second step of the LfO, to build the robot-independent task-decoder from task models to actions. For hardware-independent task-decoding, hardware-independent skill agents corresponding to each task are pre-designed and stored in the library of skill agents. These skill agents are represented only by hand motions to absorb the structural difference between each robot, and the robot arm and body is regarded as a carrier used to move the hand on the desired trajectory. Given a target trajectory by a skill agent, a general inverse kinematics (IK) solver, body role division (Sasabuchi et al. (2021)), is used to determine the robot’s body motion to achieve this trajectory while maintaining reusability for a variety of robots.
@@ 2 Related work
Efforts to increase reusability for robot programs, such as Robot Operating System (ROS) (Quigley (2009)) and OpenRTM (Ando et al. (2008)), have been conducted thus far. These two pieces of middleware follow a so-called subsumption architecture (Brooks (1986)). In this architecture, a robot program are created by combining several nodes. The nodes are properly connected and communicate with each other to execute the program. These two pieces of middleware achieved reusability by 1) unifying the format of communication and 2) providing means of communication (e.g., publisher and subscriber in ROS). Switching between low-level nodes (e.g., nodes to output sensor reading) is very easy.
@@
Recently, several papers showed the importance on the awareness of action primitives, which correspond to tasks/skill-agents in the proposed system, in the learning-from-demonstration framework. Lin et al. (2022) succeeded in the complicated tasks with multiple action primitives by training each primitive independently and combining them with sub-goals. Edmonds et al. (2019) succeeded to open medicine bottles with different locking mechanisms by learning discrete haptic states and grammatical representation of action primitives. The proposed system will prove that the goal oriented action primitives contribute to the hardware-level reusability, too.
@@ 3.1 Overview
In this paper, two terms, task and skill agent, are frequently used. A task is defined as a unit operation of what-to-do, obtained in the encoding part of LfO and on the decoding side, the procedure for a robot to perform this is referred to as a skill agent. Tasks and skill agents correspond one-to-one under an assumption of usage of one particular robot hardware, and in this paper, unless otherwise specified, they are used interchangeably without confusion. This paper aims to remove this constraint, differences of skill agents among robots, and to design skill agents that can be reused among a wide range of robot hardware. In each skill agent, the object displacement due to the hand motion is calculated to satisfy constraints from the environment. The displacement in unconstrained subspace is assumed to be obtained from the demonstration. The parameters required to perform these skill agents are called skill parameters.
In this paper, we assume that the environment between in demonstration and in robot execution is not dramatically changed. The demonstration includes a hint for collision avoidance and a robot can execute the target task by following the demonstration. Of course, we admit a slight difference in the environment in robot execution; the system can absorb the difference using sensor feedback in each skill agent.
The grasped object is in contact with the environment. A single operation of the robot, consisting of translational and rotational motions, induces contact-state transitions between the object’s surface and the environment. In a previous paper (Ikeuchi et al. (2021)), we defined the necessary and sufficient set of tasks that should be prepared as robot manipulations based on translational and rotational transitions. In this paper, we design skill agents to execute these defined tasks, assuming force feedback and visual feedback. During the design of each skill agent, we utilize changes in forces from the environment and/or visual features of the environment according to surface contact transitions for deriving reward functions that ensure successful transitions. These skills are pre-trained through reinforcement learning based on these reward functions.
Traditionally, reinforcement learning (RL) has been employed to adjust the trajectory of a robot’s hand, taking into account drag forces from the environment. Previous RL methods have predominantly focused on reward function design for specific operations, limiting their applicability and requiring individual learning for each new operation. We propose grouping multiple operations based on the types of physical constraints, deriving general guidelines, and proposing reward functions applicable to various operations based on these guidelines.
@@ 3.2.1 Surface contact and contact transition
Figure 2: Translational states and rotational states. (a) Translational states. For the sake of simplicity, we grouped three partial translational states (i.e., a hemisphere (PC1), a crescent (PC2), and a polygonal shaped state (PCN)) into one PC state, and two one-way prismatic translational states (i.e., a hemi-circle (OT1) and an arc-shaped state (OT2)) into one OT state. (b) Rotational states. See Ikeuchi et al. (2021)
@@
Figure 2: Translational states and rotational states. (a) Translational states. For the sake of simplicity, we grouped three partial translational states (i.e., a hemisphere (PC1), a crescent (PC2), and a polygonal shaped state (PCN)) into one PC state, and two one-way prismatic translational states (i.e., a hemi-circle (OT1) and an arc-shaped state (OT2)) into one OT state. (b) Rotational states. See Ikeuchi et al. (2021)
@@ 3.2.3 State transitions and skill agents
Figure 3: Translational tasks and rotational tasks.
@@ 3.3 Design principles for translational skill-agents
We will design the reward functions based on the dimensional transitions. Although transitions themselves occur through infinitesimal translation or rotation, we design skill agents by assuming actual skills performed by a robot, and that the current state persists for an finite (not infinitesimal) interval before a transition, then, the state transition occurs, and finally a new state persists for another finite interval.
When designing the reward functions for translational skill-agents, we will provide separate principles for the transition along the direction of motion and those orthogonal to the motion. This is because the transition along the direction of motion mainly affects the termination conditions of the skill agent, while directions orthogonal to the motion are related to motion control strategies during motion. We will first provide principles for the direction of motion and then consider principles for directions orthogonal to the motion.
@@ 3.3.1 Transition along the motion
Drag force that was absent in a maintenance d-state arises upon contact with an environment surface, occurring at the point of transitioning to the detachment d-state. Therefore, the occurrence of this drag force serves as the termination condition for this skill. Under normal circumstances with no errors, the contact position should align with the position given by the demonstration. However, for the sake of operational robustness to allow more errors between the demonstration and the execution, the occurrence of the drag force from the environment is considered as the termination condition for the skill agent.
@@
By moving in a admissible semi-direction in the detachment dimension, the object moves away from the environment contact surface, leading to the transition from the detachment d-state to the maintenance d-state. The disappearance of the drag force that exists in the detachment d-state can be considered as the termination condition for the skill. However, given the existence of motion in the maintenance d-state within a finite interval, the disappearance of the drag force and the the achievement of goal position are considered as the terminal conditions for this skill.
@@ 3.3.2 Transition in the dimension orthogonal to the motion
When considering transitions in the dimension orthogonal to the motion direction, nine cases can be contemplated. See Figure 5. In the following, the dimension considered is denoted as T, and the drag force encountered from the environment along this direction is represented as F-t.
Figure 5: State transition in the dimension orthogonal to the motion
@@
A moving object to have a maintenance d-state in the orthogonal direction to the motion is not constrained by the environment in that direction; therefore, the position error in that direction can be tolerated and the end position of the skill in this direction is solely determined by the demonstration.
@@
For the transition, it is necessary to adjust the motion direction to achieve surface contact in the detachment d-state. Since there is no surface contact in the orthogonal direction in the maintenance d-state, the adjustment needs to be done using positional information from visual sensors. Let us denote the position of the contact surface in the orthogonal direction as feature-t. For example, adjusting the position in the T coordinate of the manipulating object to the boundary edge of the surface that will be in contact gives an advantage in accomplishing the task. The skill aligns the positional information in the orthogonal direction using this value.
@@
In the finite interval before the transition, the motion direction is adjusted to maintain the surface contact in the orthogonal direction, i.e., detachment d-state in this direction. For this, the drag force F-t should be no greater than delta-collision so that the object dips into the environmental surface. However, an additional condition imposing non-zero drag to prevent the separation is not applied because the early departure in the orthogonal to the motion is acceptable.
After the transition, in the maintenance d-state, there are no constraints from the environment in this orthogonal direction, allowing the termination condition based on positional information given by the demonstration. However, to prevent errors due to positional inaccuracies, a condition to confirm that this direction is in the maintenance d-state is added.
@@
Generally, the transition from the detachment d-state to the constraint d-state incurs costs. Fortunately, by implementing control to retain the detachment d-state, it is possible to maintain the constraint d-state; before the transition, by retaining the detachment d-state, that is by maintaining the contact with one of the surface, the contact with the other can be maintained and after the transition, the target contact is still maintained. Therefore, the approach is adopted to maintain the detachment d-state throughout the entire interval and to keep the drag within a certain range.
@@
The constraint d-state before the transition requires to ensure that the drag force in the orthogonal direction does not exceed a threshold, delta-collision. After the transition, since the direction becomes the maintenance d-state, there is no need to check this condition. However, regardless, this condition is satisfied anyway during the maintenance d-state, there is no need to void this condition after the transition by checking the transition.
In the finite interval after the transition, the maintenance d-state allowing the use of positional information given by the demonstration. However, the condition to confirm the attainment of the maintenance d-state should be included.
@@ 3.4 Interstate transition
In this subsection, we apply the design principles obtained in the previous section to the interstate transitions of an object and derive the reward functions. Penalty conditions in the design principles are applied with OR logic, since the skill has failed if even one penalty condition is satisfied. On the other hand, the reward conditions are applied with AND logic, since the skill has reached to the goal when all the conditions are satisfied. In the design of each state transition below, we first consider the transitions from states with more constraints to states with fewer constraints and then complete the reverse transitions.
@@ 3.4.1 PC-NC and NC-PC
An object in the PC (M=2, D=1, C=0) has two maintenance d-states and one detachment d-state. For example, it could be a cube on a desk. The normal direction of the desk surface serves as the pure detachment direction (detachment d-state) and any direction along the desk surface is a maintenance direction (maintenance d-state). In the PC state, we can consider two types of skills to move the object: in the detachment direction and in the maintenance direction. See Figure 6.
PC-NC detachment skill: the pure detachment direction of the object is along the normal direction of the contact surface. The motion of the object along the direction results in no contact with surface, causing the transition to the maintenance d-state from the detachment d-state and the state of the object becomes (M=3, D=0, C=0) from (M=2, D=1, C=0). This is a typical example of the Pick-up skill.
PC-NC maintenance skill: When moving in the maintenance direction, the infinitesimal motion does not cause any transition of dimensions. The transition to NC (M=3, D=0, C=0) and loss of contact only occurs due to the shape of the environment surface. Namely, the transition occurs by moving of the object along the contact surface with maintaining contact with the surface and then reaching the edge of the surface due to the overall shape of the surface. After this point, there is no contact with the surface, and the direction orthogonal to the motion transits from the detachment d-state to the maintenance d-state.
@@
Figure 6: Two types of motion directions. (a) motion in the detachment direction. (b) motion in the maintenance direction
@@
This scenario corresponds to a typical Pick-up skill. Concerning the motion direction S, a transition occurs from the detachment d-state to the maintenance d-state, and A3 can be applied:
@@
  Reward PC-NC-a (PTG11 (Pick) skill)
@@
  B1: if T = goal-t, then reward.
On the other hand, the other dimension, the vertical direction in the example (here we call the U direction) undergoes a transition from the detachment d-state to the maintenance d-state. B5 can be applied. In the finite interval before the transition, the maintenance d-state is preserved, and in the finite interval after the transition, position control becomes relevant. However, it is necessary to include the condition to confirm that departure from the surface has occurred.
@@
        if F-u < delta-zero AND T = goal-t,
@@
      if F-u < delta-zero AND S = goal-s AND
         T = goal-t AND U = goal-u,
@@
In the same way as PC-NC, NC-PC also has two scenarios as well. Here, please imagine cases where the directions of the arrows are reversed in Figure 6. One corresponds to the common Place skill where the detachment d-state is achieved by moving from the direction that will become the detachment direction after the transition, i.e., placing an object from above the contact surface. The other is a race case where the detachment d-state is achieved by approaching the contact surface from the side, causing contact.
@@
In the two orthogonal directions to the motion, the maintenance state remains unchanged. B1 can be applied:
@@
  Reward NC-PC-a (PTG13 (Place) skill)
@@
In the example shown in the figure, a skill where the object moves parallel to the desk surface from outside of the desk to precisely induces the surface contact on the desk. This skill is not advantageous for robot operations, so its frequency of use may not be high. However, it is included here to ensure overall completeness and necessity.

### 20 Mar 2024 v2

@@ Abstract
To use new robot hardware in a new environment, it is necessary to develop a control program tailored to the specific robot in the environment. Considering the reusability of software among robots is crucial to minimize the effort involved in this process and maximize software reuse across different robots in different environments. Although recent generative AI has made it possible to automatically generate some software, robot programs are still difficult to generate because of robot’s physical interaction with the environment. This paper proposes a method to remedy this process by considering hardware-level reusability, using a Learning-from-observation (LfO) framework with a pre-designed skill-agent library. The LfO framework represents the required actions in hardware-independent representations, referred to as task models, from observing human demonstrations, capturing the necessary parameters for the interaction between the environment and the robot (Ikeuchi et al. (2021)). When executing the desired actions from the task models, a set of skill agents is employed to convert the representations into robot commands. This paper focuses on the latter part of the LfO framework, utilizing the skill-agent set to generate robot actions from the task models, and explores a hardware-independent design approach for these skill agents. These skill agents are described in a hardware-independent manner, considering the relative relationship between the robot’s hand position and the environment. As a result, it is possible to execute these actions on robots with different hardware configurations by simply swapping the inverse kinematics solver. This paper, first, defines a necessary and sufficient skill-agent set corresponding to cover all possible actions, and considers the design principles for these skill agents. We provide concrete examples of such skill agents and demonstrate the practicality of these skill agents by showing that the same representations can be executed on two different robots, Nextage and Fetch, using the proposed skill-agent set.
@@ 1 Introduction
Robot developers develop various types of robots, such as bipedal robots and mobile manipulators for satisfying users’ various demands. If we consider manipulation aspects, there are various types: a single-arm robot or a dual-arm robot, and the degrees of freedom (DOF) of an arm is from 5 (e.g., HSR, Toyota) to 7 (e.g., Fetch Mobile Manipulator, Fetch Robotics). Some robots have additional DOF on their waist.
@@
We have been developing a robot system based on a Learning-from-Observation (LfO) framework (Ikeuchi et al. (2021)) [^1] In the machine-learning (ML) community, Learning-from-Observation (LfO) is used with a slightly different definition. However, following the rationale mentioned in the introduction of Ikeuchi et al. (2021), LfO is specifically referred to as a method that transforms input into symbolic representations based on top-down knowledge and subsequently maps them to robot actions.. Our goal is to have a robot reproduce the target behavior by simply demonstrating it in front of the robot. Thus, even users without robotics knowledge can have the robot perform tasks that they desire. Unlike similar frameworks, such as learning-from-demonstration and imitation learning (Schaal (1999); Schaal et al. (2003); Billard et al. (2008); Asfour et al. (2008); Dillmann et al. (2010); Akgun et al. (2012)), the LfO system preforms indirect behavior mimicry by first a task-encoding step converting the demonstration into an abstract intermediate representation, referred to as a task model, and then a task-decoding step converting this intermediate representation into the behavior of each robot. The task-encoding recognizes behavior as symbolic task sequences (such as a sequence of pick up, place, and release actions) and then extracts from the demonstration the parameters that are pre-defined in each task (such as where to grasp and where to put). Because only task-specific information is extracted from the demonstration, unnecessary parts are ignored, allowing the demonstrator to focus only on the important demonstration parts that need to be taught. For task-decoding, we prepare in advance the skill agents, which correspond to the agents performing symbolic actions, following the design of the task model. Then, the reproduction by a robot is realized by activating the corresponding skill agent with the observed parameters.
This paper focuses on the second step of the LfO, to build the robot-independent task-decoder from task models to actions. For task-decoding, hardware-independent skill agents corresponding to each task are pre-designed and stored in the library. These skill agents represent the tasks only by hand motions to absorb the structural difference between each robot, and the robot arm and body is regarded as a carrier used to move the hand on the desired trajectory. Given a target trajectory by a skill agent, a general inverse kinematics (IK) solver, body role division (Sasabuchi et al. (2021)), is used to determine the robot’s body motion to achieve this trajectory.
@@ 2 Related work
Efforts to increase reusability of robot programs, such as Robot Operating System (ROS) (Quigley (2009)) and OpenRTM (Ando et al. (2008)), have been conducted thus far. These two pieces of middleware follow a so-called subsumption architecture (Brooks (1986)). In this architecture, a robot program are created by combining several nodes. The nodes are properly connected and communicate with each other to execute the program. These two pieces of middleware achieved reusability by 1) unifying the format of communication and 2) providing means of communication (e.g., publisher and subscriber in ROS). Switching between low-level nodes (e.g., nodes to output sensor reading) is very easy.
@@
Recently, several papers showed the importance on the awareness of action primitives, which correspond to tasks/skills in the proposed system, in the learning-from-demonstration framework. Lin et al. (2022) succeeded in the complicated tasks with multiple action primitives by training each primitive independently and combining them with sub-goals. Edmonds et al. (2019) succeeded to open medicine bottles with different locking mechanisms by learning discrete haptic states and grammatical representation of action primitives. The proposed system will prove that the goal oriented action primitives contribute to the hardware-level reusability, too.
@@ 3.1 Overview
In this paper, two terms, task and skill, are frequently used. A task is defined as a unit operation of what-to-do, obtained in the encoding part of LfO and on the decoding side, the procedure for a robot to perform this is referred to as a skill and the skill agent are responsible for the execution of a skill. Tasks and skills correspond one-to-one under an assumption of usage of one particular robot hardware, and in this paper, unless otherwise specified, they are used interchangeably without confusion. This paper aims to remove this constraint, i.e., differences of skills among robots, and to design skill agents that can be reused among a wide range of robot hardware. In each skill agent, the object displacement due to the hand motion is calculated to satisfy constraints from the environment. The displacement in unconstrained subspace is assumed to be obtained from the demonstration. The parameters required to perform these skills are called skill parameters.
In this paper, we assume that the environment between in demonstration and in robot execution is not dramatically changed. The demonstration includes a hint for collision avoidance and a robot can execute the target task by following the demonstration. Of course, we admit a slight difference in the environment in robot execution; the skill agent can absorb the difference using sensor feedback.
The grasped object is in contact with the environment. A single operation of the robot, consisting of translational and rotational motions, induces contact-state transitions between the object’s surface and the environment. In a previous paper (Ikeuchi et al. (2021)), we defined the necessary and sufficient set of tasks that should be prepared as robot manipulations based on translational and rotational transitions. In this paper, we design skill agents to execute these defined tasks, assuming force feedback and visual feedback. During the design of each skill agent, we utilize changes in forces from the environment and/or visual features of the environment according to surface contact transitions and derive reward functions that ensure successful transitions using them. These skill agents are pre-trained through reinforcement learning based on these reward functions.
Traditionally, reinforcement learning (RL) has been employed to adjust the trajectory of a robot’s hand, taking into account drag forces from the environment. Previous RL methods have predominantly focused on the design of reward functions for specific operations and requiring individual learning for each new operation coming. We propose grouping multiple operations based on the types of physical constraints, deriving general guidelines, and proposing reward functions applicable to various operations based on these guidelines.
@@ 3.2.1 Surface contact and contact transition
Figure 2: Translational states and rotational states. (a) Translational states. For the sake of simplicity, we grouped three partial translational states (i.e., a hemisphere (PC1), a crescent (PC2), and a polygonal shaped state (PCN)) into one PC state, and two one-way prismatic translational states (i.e., a hemi-circle (OT1) and an arc-shaped state (OT2)) into one OT state. (b) Rotational states. See Ikeuchi et al. (2021).
@@ 3.2.3 State transitions and skill agents
Figure 3: Translational tasks and rotational tasks
@@ 3.3 Design principles for translational skill-agents
We will design the reward functions based on the dimensional transitions. Although transitions themselves occur through infinitesimal translation or rotation, we design skill agents by assuming actual skills performed by a robot, that is, the current state persists for an finite (not infinitesimal) interval before a transition, then, the state transition occurs, and finally a new state persists for another finite interval.
When designing the reward functions for translational skills, we will provide separate principles for the transition along the direction of motion and those orthogonal to the motion. This is because the transition along the direction of motion mainly affects the termination conditions of the skill agent, while directions orthogonal to the motion are related to motion control strategies during motion. We will first provide principles for the direction of motion and then consider principles for directions orthogonal to the motion.
@@ 3.3.1 Transition along the motion
Drag force that was absent in a maintenance d-state arises upon contact with an environment surface, occurring at the point of transitioning to the detachment d-state. Therefore, the occurrence of this drag force serves as the termination condition for this skill. Under normal circumstances with no errors, the contact position should align with the position given by the demonstration. However, for the sake of operational robustness to allow more gaps between the demonstration and the execution, the occurrence of the drag force from the environment is considered as the termination condition for the skill agent.
@@
By moving in a admissible semi-direction in the detachment dimension, the object moves away from the environment contact surface, leading to the transition from the detachment d-state to the maintenance d-state. The disappearance of the drag force that exists in the detachment d-state can be considered as the termination condition for the skill. However, given the existence of motion in the maintenance d-state within a finite interval, the disappearance of the drag force and the the achievement of goal position are considered as the terminal conditions for this skill agent.
@@ 3.3.2 Transition in the dimension orthogonal to the motion
When considering transitions in the dimension orthogonal to the motion direction, nine cases can occur. See Figure 5. In the following, the dimension considered is denoted as T, and the drag force encountered from the environment along this direction is represented as F-t.
Figure 5: State transitions in the dimension orthogonal to the motion
@@
A moving object to have a maintenance d-state in the orthogonal direction to the motion is not constrained by the environment in that direction. Therefore, the position error in that direction can be tolerated and the end position of the skill in this direction is solely determined by the demonstration.
@@
For the transition, it is necessary to adjust the motion direction to achieve surface contact in the detachment d-state. Since there is no surface contact in the orthogonal direction in the maintenance d-state, the adjustment needs to be done using positional information from visual sensors. Let us denote the position of the contact surface in the orthogonal direction as feature-t. For example, adjusting the position in the T coordinate of the manipulating object to the boundary edge of the surface that will be in contact gives an advantage in accomplishing the task. The skill agent aligns the positional information in the orthogonal direction using this value.
@@
In the finite interval before the transition, the motion direction is adjusted to maintain the surface contact in the orthogonal direction, i.e., detachment d-state in this direction. For this, the drag force F-t should be no greater than delta-collision so that the object dips into the environmental surface. In strict sense, the detachment d-state should be kept until the surface contact is disappeared to the motion (i.e., detach it from the boundary as shown in Figure 6 (b)). Then an additional condition imposing non-zero drag to prevent the separation is applied.
After the transition, in the maintenance d-state, there are no constraints from the environment in this orthogonal direction, allowing the termination condition based on positional information given by the demonstration.
@@
Generally, the transition from the detachment d-state to the constraint d-state incurs costs. Fortunately, by implementing control to retain the detachment d-state, it is possible to achieve the constraint d-state. Before the transition, retaining the detachment d-state is realized by maintaining the contact with one of the surface. And the contact with the other is automatically achieved after the transition because of the geometries of the object and environment, and the original contact is still maintained. Therefore, the skill agent maintains the detachment d-state throughout the entire interval and to keep the drag within a certain range.
@@
The constraint d-state before the transition requires to ensure that the drag force in the orthogonal direction does not exceed a threshold, delta-collision. After the transition, since the direction becomes the maintenance d-state, there is no need to check this condition. In the finite interval after the transition, the maintenance d-state allowing the use of positional information given by the demonstration. However, the condition to confirm the attainment of the maintenance d-state should be included.
@@ 3.4 Interstate transition
In this subsection, we apply the design principles obtained in the previous section to the interstate transitions of an object and derive the reward functions. Penalty conditions in the design principles are applied with OR logic, since the task has failed if even one penalty condition is satisfied. On the other hand, the reward conditions are applied with AND logic, since the task has reached to the goal when all the conditions are satisfied. In the design of each state transition below, we first consider the transitions from states with more constraints to states with fewer constraints and then complete the reverse transitions.
@@ 3.4.1 PC-NC and NC-PC
An object in PC (M=2, D=1, C=0) has two maintenance d-states and one detachment d-state. For example, it could be a cube on a desk. The normal direction of the desk surface serves as the pure detachment direction (detachment d-state) and any direction along the desk surface is a maintenance direction (maintenance d-state). In PC, we can consider two types of tasks to move the object: in the detachment direction and in the maintenance direction as shown in Figure 6.
PC-NC detachment motion: the pure detachment direction of the object is along the normal direction of the contact surface. The motion of the object along the direction results in no contact with surface, causing the transition to the maintenance d-state from the detachment d-state and the state of the object becomes NC (M=3, D=0, C=0) from PC. This is a typical example of the Pick-up task.
PC-NC maintenance motion: When moving in the maintenance direction, the infinitesimal motion does not cause any transition of dimensions. The transition to NC and loss of contact only occurs due to the shape of the environment surface. Namely, the transition occurs by moving of the object along the contact surface with maintaining contact with the surface and then reaching the edge of the surface due to the overall shape of the surface. After this point, there is no contact with the surface, and the direction orthogonal to the motion transits from the detachment d-state to the maintenance d-state.
@@
Figure 6: Two types of motion directions. (a) motion in the detachment direction. (b) motion in the maintenance direction.
@@
This scenario corresponds to a typical Pick-up task. Concerning the motion direction S, a transition occurs from the detachment d-state to the maintenance d-state, and A3 can be applied:
@@
  Reward PC-NC-a (PTG11 (Pick) task)
@@
  B1: if T = goal-t, then reward
On the other hand, the other dimension, the vertical direction in the example (here we call the U direction) undergoes a transition from the detachment d-state to the maintenance d-state. B5 can be applied. In the finite interval before the transition, the detachment d-state is preserved, and in the finite interval after the transition, position control becomes relevant. However, it is necessary to include the condition to confirm that departure from the surface has occurred.
@@
        if F-u < delta-zero AND U = goal-u,
@@
      if S = goal-s AND T = goal-t AND
         F-u < delta-zero AND U = goal-u,
@@
In the same way as PC-NC, NC-PC also has two scenarios. Here, please imagine cases where the directions of the arrows are reversed in Figure 6. One corresponds to the common Place task where the detachment d-state is achieved by moving from the direction that will become the detachment direction after the transition, i.e., placing an object from above the contact surface. The other is a race case where the detachment d-state is achieved by approaching the contact surface from the side, causing contact.
@@
In the two orthogonal directions to the motion, the maintenance d-state remains unchanged. B1 can be applied:
@@
  Reward NC-PC-a (PTG13 (Place) task)
@@
In the example shown in the figure, the object moves parallel to the desk surface from the outside of the desk to precisely induces the surface contact on the desk. This task is not advantageous for robot operations, so its frequency of use may not be high. However, it is included here to ensure overall completeness and necessity.